the speedometer of a motor car showed the following speeds in km/hr at the ends of successive intervals of 3 secs: 38.9, 52.8, 64.4, 73.6, 80.9, 86.1, 90.0, 92.5 plot the velocity time graph in convenient units and find from the graph the distance covered in metre while the speed increased from 80 to 90kmph, and the accelerations in SI units when the speeds are 50, 70, 90kmph respectively


1
Expert's answer
2021-08-09T06:51:10-0400
"s=158\ m"


"a_1=1.5\frac{m}{s^2}\\a_2=0.73\frac{m}{s^2}\\a_3=0.35\frac{m}{s^2}\\"